# High-Temperature Fermionic Gibbs States are Mixtures of Gaussian States

### v1

@@ 5.3 Sample Tree
for some $A$. The second summand is at most $\log(25)$, since in any step $\sigma_{v}$ is multiplied by a commuting operator with trace 1 or trace $1\pm\alpha_{v}$. This latter value lies in $[\frac{1}{25},1+\frac{24}{25}]$, and therefore the logarithm is bounded in absolute value by $\log(25)$. Since $\beta\leq\frac{1}{50\mathcal{R}d^{2}}\leq\frac{1}{100d}$, we can take $C=25$, so $\lVert A\rVert_{\infty}\leq\frac{1}{49}$. Lemma 5.6 applies again since $e^{-\beta HS_{w}}$ is positive, yielding an upper bound of $\log(25\cdot 49\cdot 48^{-1})$, including the second summand.
@@ 5.4 Truncated Sample Tree
Specifically, $A$ is the event that $t_{k}\leq t_{\text{max}}$ for threshold $t_{\text{max}}=\lceil\log_{2}\left(\delta^{-1}\right)\rceil$. Moreover, the distribution $\mu^{\prime}$ has support with size $\text{exp}(\text{polylog}\left(\delta^{-1}\right))$.
@@
This algorithm to perform this sampling will proceed in two steps. Firstly, it samples a value of $t=1,\dots,t_{\text{max}}$ from the distribution $\frac{2^{-t}}{1-2^{-t_{\text{max}}}}$, which may be done exactly and efficiently. Then, given the sampled $t$, it must sample $I$ with probability $1-c_{t}(2N\mathcal{R})^{t}$ and each $I+(N\mathcal{R})^{-t}\Gamma_{j,t}$ with probability $c_{j,t}(2N\mathcal{R})^{t}$. However, Corollary 3.3 is able to sample each $\Gamma_{j,t}$ with probabilities $p_{j,t}\geq c_{j,t}(2N\mathcal{R})^{t}$, in time. After sampling an operator $\Gamma_{j,t}$, the algorithm can further sample from $\{0,\Gamma_{j,t}\}$ with probabilities $1-\frac{p_{j,t}}{(2N\mathcal{R})^{-t}}$ and $\frac{p_{j,t}}{c_{j,t}(2N\mathcal{R})^{-t}}$. We are now sampling each $\Gamma_{j,t}$ with probability $c_{j,t}(2N\mathcal{R})^{-t}$ and 00 with the remaining probability. Returning $I+(2N\mathcal{R})^{-t}\Gamma_{j,t}$ or $I$ correspondingly gives the desired algorithm.
@@
6 Conclusion
Motivated by recent progress on the lack of entanglement in high-temperature Gibbs states of quantum spin systems [5], in this work we showed that high-temperature Gibbs states of fermionic systems are probabilistic mixtures of fermionic Gaussian states. Fermionic Gaussian states are completely determined by their correlation matrices, which enables efficient classical representation of these states. Based on this structural result, we then proposed an efficient algorithm to sample from this distribution of fermionic Gaussian states, thus enabling efficient computation of observable expectation values and sampling from the Gibbs states.
Besides techniques developed in [5], our results rely on a new adaptive pairing procedure for Majorana modes. While for quantum spin systems locality of the operators can imply commutativity, such connection is only possible in the fermionic case with an appropriate pairing.
Our result sharpens the the boundary between quantum advantage and classical simulability for fermionic systems. Recent results on the mixing time of weakly interacting fermionic systems shows that at any constant temperature, the corresponding Gibbs states can be efficiently prepared on a quantum computer [20, 21]. Our result shows that in the high-temperature regime, this mixing time result does not lead to a super-polynomial quantum advantage. However, for a temperature below the threshold considered in this work, the existence of significant quantum advantage remains an open problem. Given the importance of the structural result in the design of the algorithm, it is natural to ask whether weakly interacting fermionic Gibbs states can also be expressed as mixtures of fermionic Gaussian states.
The connection between fermionic Gaussian mixtures and simulability was also explored in the simulation of fermionic circuits [26]. Given the prominent role played by this structure in the simulability of fermionic systems, it is of interest to explore ways to test whether a fermionic quantum state is a mixture of Gaussian states or not. This question can be asked in at least two different settings: Firstly, given a quantum state that is the ground state or the Gibbs state of a non-quadratic fermionic Hamiltonian, can we distinguish whether it is a mixture of fermionic Gaussian states? Secondly, if we have an unknown quantum state on an experimental fermionic system, what are the sample and computational complexities of answering such a question? Answering these questions can help us better understand the classical simulability of fermionic systems and the boundary of quantum advantage.
@@ Acknowledgments
We thank Ewin Tang, Ainesh Bakshi, and Robbie King for helpful discussions. Y.C. is supported by the Mellon Mays Undergraduate Fellowship. J.J. is supported by MURI Grant FA9550-18-1-0161, NSF CAREER award CCF-2048204, and the IQIM, an NSF Physics Frontiers Center (NSF Grant PHY-1125565).

### v2

@@ 5.3 Sample Tree
for some $A$. The second summand is at most $\log(25)$, since in any step $\sigma_{v}$ is multiplied by a commuting operator with trace 1 or trace $1\pm\alpha_{v}$. This latter value lies in $\frac{1}{25},1+\frac{24}{25}]$, and therefore the logarithm is bounded in absolute value by $\log(25)$. Since $\beta\leq\frac{1}{50\mathcal{R}d^{2}}\leq\frac{1}{100d}$, we can take $C=25$, so $\lVert A\rVert_{\infty}\leq\frac{1}{49}$. Lemma 5.6 applies again since $e^{-\beta HS_{w}}$ is positive, yielding an upper bound of $\log(25\cdot 49\cdot 48^{-1})$, including the second summand.
@@ 5.4 Truncated Sample Tree
Specifically, AA is the event that $t_{k}\leq t_{\text{max}}$ for threshold $t_{\text{max}}=\lceil\log_{2}\left(\delta^{-1}\right)\rceil$. Moreover, the distribution $\mu^{\prime}$ has support with size $\text{exp}(\text{polylog}\left(\delta^{-1}\right))$.
@@
This algorithm to perform this sampling will proceed in two steps. Firstly, it samples a value of $t=1,\dots,t_{\text{max}}$ from the distribution $\frac{2^{-t}}{1-2^{-t_{\text{max}}}}$, which may be done exactly and efficiently. Then, given the sampled $t$, it must sample $I$ with probability $1-c_{t}(2N\mathcal{R})^{t}$ and each $I+(N\mathcal{R})^{-t}\Gamma_{j,t}$ with probability $c_{j,t}(2N\mathcal{R})^{t}$. However, Corollary 3.3 is able to sample each $\Gamma_{j,t}$ with probabilities $p_{j,t}\geq c_{j,t}(2N\mathcal{R})^{t}$, in time. After sampling an operator $\Gamma_{j,t}$, the algorithm can further sample from $\{0,\Gamma_{j,t}\}$ with probabilities $1-\frac{p_{j,t}}{(2N\mathcal{R})^{-t}}$ and $\frac{p_{j,t}}{c_{j,t}(2N\mathcal{R})^{-t}}$. We are now sampling each $\Gamma_{j,t}$ with probability $c_{j,t}(2N\mathcal{R})^{-t}$ and 0 with the remaining probability. Returning $I+(2N\mathcal{R})^{-t}\Gamma_{j,t}$ or $I$ correspondingly gives the desired algorithm.
@@
6 High-Temperature SYK is Non-Gaussian
In this section, we present an example of fermionic Hamiltonians whose Gibbs states at high temperature (small $\beta$) cannot be expressed as a probabilistic mixture of fermionic Gaussian states. Specifically, we focus on the Sachdev–Ye–Kitaev (SYK) model [33, 21]. The $q$-local SYK model for even $q$ is defined as
$H=i^{q/2}\sum_{1\leq k_{1}<...<k_{q}\leq 2n}J_{k_{1}\dots k_{q}}\gamma_{k_{1}}\dots\gamma_{k_{q}}.$ (60)
where $\{J_{k_{1}\dots k_{q}}\}_{k_{1}\dots k_{q}}$ are independent Gaussian variables with mean 0 and variance 1. We use $H\sim\text{SYK}_{q}$ to denote sample a random $\text{SYK}_{q}$ Hamiltonian according to the randomness of $\{J_{k_{1}\dots k_{q}}\}_{k_{1}\dots k_{q}}$.
It was shown that Gaussian states cannot achieve constant approximation ratio for the ground state to the SYK model [15, 17, 13]. In this section, we show that those results can be generalized to Gibbs states of the SYK model at constant temperature, formally proven in Theorem 6.5.
In particular, $\text{SYK}_{q}$ consists of constant locality Hamiltonians, but not of constant degree, so our structural result (Theorem 4.2) does not apply. Theorem 6.5 shows that a large class of random Hamiltonians of constant locality do not exhibit death of entanglement.
Here, we use $negl(n)$ to denote negligible functions which decay exponentially fast with $n$ and $||\cdot||$ to denote operator norms. Furthermore, we use the asymptotic notation $\omega$ so that $f=\omega(g)$ means that $g=o(f)$, i.e. $\frac{g}{f}$ asymptotically vanishes.
Lemma 6.1
There exists a constant $c>0$ such that with probability 1-$\text{negl}(n)$,
$\lVert H\rVert\leq cn^{(q+1)/2},$ (61)
for some sufficiently large constant $c>0.$ In other words, with all but negligible probability over $H\sim\text{SYK}_{q}$, $\lVert H\rVert=O(n^{(q+1)/2})$.
Proof.
We apply the Gaussian matrix Chernoff bound to
$H=i^{q/2}\sum_{1\leq k_{1}<\dots<k_{q}\leq 2n}J_{k_{1}\dots k_{q}}\gamma_{k_{1}}\dots\gamma_{k_{q}},$ (62)
which applies since $J_{k_{1}\dots k_{q}}$ is a standard normal random variable and $i^{q/2}\gamma_{i_{1}}\dots\gamma_{i_{q}}$ is Hermitian. Applying the bound directly, we obtain
$\mathbb{P}\left[\lVert H\rVert>t\right]$ $\leq 2^{n}\exp\left(-\frac{t^{2}}{2\sigma^{2}}\right),$ (63)
where
$\sigma^{2}=\left\|\sum_{1\leq k_{1}<\dots<k_{q}\leq 2n}(i^{q/2}\gamma_{k_{1}}\dots\gamma_{k_{q}})^{2}\right\|_{2}={2n\choose q}$ (64)
where $\|\cdot\|_{2}$ is the matrix $2$-norm.
Since ${2n\choose q}\leq Dn^{q}$ for sufficiently large constant $D>0$. Choosing $t=cn^{(q+1)/2}$ for sufficiently large constant $c>0$ such that $c^{2}>2\ln(2)D$, we find that $\mathbb{P}\left[\lVert H\rVert>t\right]\leq 2^{n}\exp\left(-\frac{c^{2}}{2D}n\right)=\text{negl}(n)$.
Lemma 6.2
For any $H\sim\text{SYK}_{q}$, $\text{tr}(H)=0$.
Proof.
For any Majorana string $\gamma_{i_{1}}\dots\gamma_{i_{k}}$ which is not equal to a scalar, $\text{tr}(\gamma_{i_{1}}\dots\gamma_{i_{k}})=0$. Since the $\text{SYK}_{q}$ distribution consists of Hamiltonians that are sums of nontrivial Majorana strings, we may conclude that $\text{tr}(H)=0$. ∎
Lemma 6.3
There exists constants $0<c_{1}<c_{2}$ such that with probability $1-\text{negl}(n)$,
$c_{1}n^{q}\leq\frac{\text{tr}(H^{2})}{2^{n}}\leq c_{2}n^{q}.$ (65)
In other words, with all but negligible probability over $H\sim\text{SYK}_{q}$, $\frac{\text{tr}(H^{2})}{2^{n}}=\Theta(n^{q})$.
Proof.
The Hamiltonian $H$ is of the form
$H=i^{q/2}\sum_{1\leq k_{1}<\dots<k_{q}\leq 2n}J_{k_{1}\dots k_{q}}\gamma_{k_{1}}\dots\gamma_{k_{q}}.$ (66)
Aftering squaring the Hamiltonian and expanding, the scalar term of $H^{2}$ is
$\sum_{1\leq k_{1}<\dots<k_{q}\leq 2n}J_{k_{1}\dots k_{q}}^{2}I.$ (67)
whose trace equals to $\text{tr}(H^{2})$, since the trace of non-trivial Majorona string is 0. Since the trace of $I$ is $\text{tr}(I)=2^{n}$, we conclude that
$\frac{\text{tr}(H^{2})}{2^{n}}=\sum_{1\leq k_{1}<\dots<k_{q}\leq 2n}J_{k_{1}\dots k_{q}}^{2}.$ (68)
Note that the value $\sum_{1\leq k_{1}<\dots<k_{q}\leq 2n}J^{2}_{k_{1}\dots k_{q}}$ follows the Chi-Square distribution thus exhibits strong concentration around its mean (the Laurent-Massart bound). Thus with probability $1-\text{negl}(n)$, this sum is $\Omega(n^{q})$. ∎
Finally, we mention a result that bounds the maximum energy that any Gaussian state can attain with high probability over the SYK distribution.
Lemma 6.4 (Lemma 9, [17])
There exists a constant $C>0$ for which with probability $1-\text{negl}(n)$,
$\left|\text{tr}(\tau H)\right|\leq Cn^{q/4+1}.$ (69)
In other words, with all but negligible probability over $H\sim\text{SYK}_{q}$ and for all Gaussian states $\tau$, $\left|\text{tr}(\tau H)\right|=O(n^{q/4+1})$.
Theorem 6.5
For any constant $\epsilon>0$ and inverse temperature $\beta=\Omega(n^{-3q/4+1+\epsilon}),$ with high probability over $H\sim\text{SYK}_{q}$ there exists some $\beta_{0}<\beta$ for which
$\rho_{\beta_{0}}=\frac{e^{-\beta_{0}H}}{\text{tr}\left(e^{-\beta_{0}H}\right)}.$ (70)
is not a convex combination of Gaussian states.
Proof.
We begin with the notation
$\langle-H\rangle_{\beta}=\text{tr}(-\rho_{\beta}H).$ (71)
By explicit calculation, we see that
$\frac{d}{d\beta}\langle-H\rangle_{\beta}=\langle H^{2}\rangle_{\beta}-\langle H\rangle_{\beta}^{2}$ (72)
and
$\frac{d^{2}}{d^{2}\beta}\langle-H\rangle_{\beta}=-\langle H^{3}\rangle_{\beta}+3\langle H\rangle_{\beta}\langle H^{2}\rangle_{\beta}-2\langle H\rangle_{\beta}^{3}.$ (73)
Note that
$D$ $\coloneqq\left(\frac{d}{d\beta}\langle-H\rangle_{\beta}\right)\Big|_{0}$ (74)
$=\frac{\text{tr}(H^{2})}{2^{n}}-\left(\frac{\text{tr}(H)}{2^{n}}\right)^{2}$ (75)
$=\Theta(n^{q}),$ (76)
where the last equality holds by Lemma 6.2 and Lemma 6.3 with all but negligible probability. Meanwhile, we may upper bound the second derivative at any $\beta$ with
$\left|\frac{d^{2}}{d^{2}\beta}\langle-H\rangle_{\beta}\right|$ $=O(\lVert H\rVert^{3})$ (77)
$=O(n^{3(q+1)/2})$ (78)
since in general $\langle H\rangle_{\beta}=\text{tr}(\rho_{\beta}H)\leq\lVert H\rVert$. The second inequality holds by Lemma 6.1 with all but negligible probability over $\text{SYK}_{q}$.
Let $S=\Theta(\lVert H\rVert^{3})$ be the upper bound on the second derivative, and let $0<\beta_{0}\leq\frac{D}{S}$. Note that for any $0\leq\beta\leq\beta_{0}$, according to Eq. (73) and Eq. (76), we have that
$\frac{d\langle-H\rangle_{\beta}}{d\beta}\geq D-\beta S.$ (79)
Then
$\langle-H\rangle_{\beta_{0}}$ $\geq\int_{0}^{\beta_{0}}(D-\beta S)d\beta$ (80)
$=D\beta_{0}-\frac{S\beta_{0}^{2}}{2}$ (81)
$\geq D\beta_{0}-\frac{D\beta_{0}}{2}$ (82)
$=\frac{D\beta_{0}}{2}.$ (83)
where the first inequality holds because $D-\beta S$ is a lower bound on the derivative of $\langle-H\rangle_{\beta}$ at $\beta$, and $\langle-H\rangle_{\beta_{0}}\left|{}_{\beta_{0}=0}\right.=tr(H)=0$. We conclude that with all but negligible probability over $H\sim\text{SYK}_{q}$
$\text{tr}(\rho_{\beta_{0}}(-H))$ $=\Omega\left(D\beta_{0}\right)=\Omega\left(n^{q}\beta_{0}\right).$ (84)
The maximum value given for $\beta_{0}$ is
$D/S$ $=\Omega(n^{q-3(q+1)/2})$ (85)
$=\Omega(n^{-q/2-3/2})$ (86)
$=\omega(n^{-3q/4+1}),$ (87)
where the last bound follows when $q\geq 12$. Hence, we may at least choose $\beta_{0}=\omega(n^{-3q/4+1})$. We now show that for this $\beta_{0}$, $\rho_{\beta_{0}}$ achieves an energy not asymptotically achievable by any Gaussian state, so as a consequence $\rho_{\beta_{0}}$ cannot be a convex combination of Gaussian states. By Lemma 6.4, the best energy a Gaussian state $\tau$ can achieve is
$|\text{tr}(\tau H)|=O(n^{q/4+1}).$ (88)
Since $D/S=\omega(n^{-3q/4+1})$, we may choose $\beta_{0}=\omega(n^{-3q/4+1})$, obtaining
$\text{tr}(\rho_{\beta_{0}}(-H))$ $=\Omega(n^{q}\beta_{0})$ (89)
$=\omega(n^{q/4+1})$ (90)
$=\omega(|\text{tr}(\tau H)|)$ (91)
$=\omega(\text{tr}(\tau(-H)))$ (92)
We conclude that with all but negligible probability over $H\sim\text{SYK}_{q}$, for any constant $\epsilon>0$ and $\beta=\Omega(n^{-3q/4+1+\epsilon})$, $\rho_{\beta_{0}}$ is not a distribution of Gaussian states. Since $\beta_{0}$ can be chosen to be asymptotically smaller than $\beta$ this proves the result.
Corollary 6.6
For typical normalizations of the SYK model, such as $\frac{1}{n^{q/2}}H$ and $\frac{1}{n^{(q-1)/2}}H$ for $H\sim\text{SYK}_{q}$, there is no constant $\beta$ such that for any $\beta^{\prime}<\beta$, the Gibbs state $\rho_{\beta^{\prime}}$ is a convex combination of Gaussian states.
Proof.
This statement follows from Theorem 6.5. In the theorem, for any $\beta=\omega(n^{-3q/4+1+\epsilon})$, we have that
$\frac{e^{-\beta H}}{\text{tr}(e^{-\beta H})}$ (93)
is not a convex combination of Gaussian states.
Dividing $H$ by the normalization factor $C=n^{q/2}$ or $C=n^{(q-1)/2}$, the same result holds when $\beta$ is scaled up by the same normalization factor. The bound on $\beta$ in the condition of the Theorem 6.5 becomes $\Omega(n^{-q/4+1+\epsilon})$ or $\Omega(n^{-q/4+1/2+\epsilon})$ in either case. For $q\geq 12$, and choosing $\epsilon=1$, any constant $\beta$ satisfies this bound, and as a consequence there is some $\beta_{0}<\beta$ for which
$\frac{e^{-\beta_{0}(H/C)}}{\text{tr}(e^{-\beta_{0}(H/C)})}$ (94)
is not a distribution of Gaussian states. ∎
@@ Acknowledgments
We thank Ewin Tang, Ainesh Bakshi, Robbie King and Eric Anschuetz for helpful discussions. Y.C. is supported by the Mellon Mays Undergraduate Fellowship. J.J. is supported by NSF CCF-2321079, MURI Grant FA9550-18-1-0161, NSF CAREER award CCF-2048204, and the IQIM, an NSF Physics Frontiers Center (NSF Grant PHY-1125565).